You want to invest an amount of money today and receive back twice that amount in the future. You expect to earn 9 percent interest. Approximately how long must you wait for your investment to double in value?

Please provide an Office Excel formula in your answer
Answer #1

Ans 8.04 years

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
Let amount of Investment be $ 100
FV / PV = (1 + r )^n
200 / 100 = (1 + 9%)^n
2 = 1.09^n
n = 8.04 years